OpenStudy (anonymous):

Choose the correct graph of the given system of equations. y + 2x = −1 3y − x = 4

OpenStudy (anonymous):

ur in 7th grade

OpenStudy (anonymous):

or none of the above.

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

@Mertsj @ash2326 @hartnn

OpenStudy (anonymous):

y + 2x = 1 sub in 0 for x y + 2(0) = 1 y + 0 = 1 y = 1 <-- y axis now sub 0 in for y 0 +2x = 1 2x = 1 x = 1/2 < -- x axis are you sure it isn't none f the above @mike1ney

OpenStudy (anonymous):